OpenStudy (anonymous):

A man rides at the rate of 18 km/hr but stops for 6 mins to change horses at the end of every 7th km .the time that he will take to cover a distance of90km is

OpenStudy (anonymous):

total time taken = 90/18=5hrs no of times horse changed=90/7=12 time taken in changing hourses = 12X6=72min=72/60hrs=1+(12/60) total time taken =5+1+0.2=6.2hrs

OpenStudy (anonymous):

have u got it?

OpenStudy (anonymous):

s.i got it.but some changes in .2 hrs

OpenStudy (anonymous):

then what is correction?

OpenStudy (anonymous):

12.85into 6 =77.14

OpenStudy (anonymous):

why 12.85?

OpenStudy (anonymous):

90/7

OpenStudy (anonymous):

k.thanks

OpenStudy (anonymous):

no no dont misconcept it its important

OpenStudy (anonymous):

be logical no of times horse changed willbe only in whole no. not in fractions

OpenStudy (anonymous):

i dont know that greater integer function .

OpenStudy (anonymous):

so its 12 not 12.85

OpenStudy (anonymous):

leave greatest integer function try to understand what i said above

OpenStudy (anonymous):

k .i have options for this question but this answer doesnt mach.a) 6hrs b) 6hrs.12min c)6 hrs 18min d) 6hrs 24min

OpenStudy (anonymous):

is answer given or not?

OpenStudy (anonymous):

given .....

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

thanks frnd.

OpenStudy (anonymous):

ok u r welcome